A 6.8 KΩ resistor has burned out in a circuit.you must replace it with another resistor with the same resistance value.if the resistor carries 10 mA what should its power rating be..?assume that you have available resistors in all the standard power ratings.


1
Expert's answer
2021-04-06T06:33:54-0400

Power "P=I^2R=(0.01)^2\times 6800=0.68" Watt


Hence The power rating must be 0.68watt

That's the bare minimum; the closest standard resistor would be 1W